### v1

# UTMath: Math Evaluation with Unit Test via Reasoning-to-Coding Thoughts

Abstract
The evaluation of mathematical reasoning capabilities is essential for advancing Artificial General Intelligence (AGI). While Large Language Models (LLMs) have shown impressive performance in solving mathematical problems, existing benchmarks such as GSM8K and MATH present limitations, including narrow problem definitions with specific numbers and reliance on predetermined rules that hinder accurate assessments of reasoning and adaptability. This paper introduces the UTMath Benchmark, which robustly evaluates the models through extensive unit tests. It consists of 1,053 problems across 9 mathematical domains, with over 68 test cases per problem. We propose an innovative evaluation framework inspired by unit testing in software development, focusing on both accuracy and reliability of results. Furthermore, we introduce the Reasoning-to-Coding of Thoughts (RCoT) approach, which encourages LLMs to perform explicit reasoning before generating code, leading to generating more advanced solution and improved performance. Furthermore, we are releasing not only the UTMath benchmark but also the UTMath-Train training dataset (more than 70k samples), to support the community in further exploring mathematical reasoning.
UTMath: Math Evaluation with Unit Test
via Reasoning-to-Coding Thoughts
1 Introduction
The pursuit of AGI necessitates strong mathematical reasoning capabilities, making the evaluation of such abilities a crucial area of research Zhou et al. (2024a). Recent advancements in LLMs have demonstrated remarkable proficiency in solving complex mathematical problems, achieving amazing performance on various datasets of Math Word Problems (MWPs), such as GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021), TheoremQA Chen et al. (2023).
Figure 1: Comparing UTMath with MATH. In the left, GPT-4o easily solved a question but failed on a similar question with a different number. The right figure illustrates our benchmark, where each problem includes multiple test cases, and a solution must pass all test cases to be strictly considered correct. We also propose a new prompting method RCoT, where the LLM first reasons through the problem and then generates code.
However, classic benchmarks exhibit several limitations that impede the accurate and comprehensive assessment of these models’ capabilities Ahn et al. (2024). First, these benchmarks test models on narrowly defined problems with some specific numbers, which do not adequately assess adaptability to similar but varied scenarios as shown in Fig. 1. Second, their evaluation relies on predetermined rules or the method of LLM-as-a-Judge( Dubois et al. (2024); Zheng et al. (2023)) that usually failed with capricious responses of LLMs. For example, an accurate answer need to be extracted to exactly match the fianl answer in the dataset GSM8K, TheoremQA, and MATH dataset. Furthermore, these benchmarks focus more on final answers than on the underlying reasoning steps. While recent work has made great progress in developing new benchmarks for evaluating the mathematical reasoning of LLMs, many of these approaches still fall short of addressing the fundamental limitations of earlier datasets. For instance, benchmarks like GSM-HARD Gao et al. (2023), GSM-IC Shi et al. (2023), GSM-Plus Li et al. (2024a), MetaMath Yu et al. (2023) have extended the dataset of GSM8K or MATH with some perturbation such as substitution, reversing, distractor insertion. These efforts, while valuable, are characterized by limited coverage and high costs.
In this context, our work seeks to bridge these gaps by proposing a solid and robust benchmark that accurately evaluates the mathematical capabilities of LLMs.Drawing inspiration from evaluation methods in software development, where solution correctness is determined through comprehensive unit testing, a solution is deemed correct if it passes all tests, as this demonstrates that its logic is sufficiently robust to generalize across a broad spectrum of scenarios. Similarly, we propose designing a comprehensive set of unit tests for mathematical problems to rigorously assess the reasoning processes of LLMs. If a solution passes all unit tests across a group of similar problems, it suggests that the reasoning behind the solution is more reliable and trustworthy. Specifically, we introduce the UTMath, a novel benchmark derived from the On-Line Encyclopedia of Integer Sequences (OEIS) OEIS Foundation Inc. (2024). The benchmark consists of 1,053 problems spanning 9 mathematical domains, such as Number Theory and Geometry. Each problem is accompanied by more than 68 test cases that provide a set of inputs and their corresponding outputs.
In terms of evaluation methodology, our benchmark requires models to derive a general solution for a class of problems, typically represented in the form of code. Compared to solving a problem defined by specific numbers, developing such a general solution is substantially more challenging, requiring higher levels of intelligence and reasoning ability. However, we observed that when the model is required to perform reasoning and coding in a single response, as in “Program of Thoughts (PoT)” Chen et al. (2022), it consistently tends to produce simpler, more straightforward solutions. We surmise that this tendency may be influenced by the distribution of coding data. To address this, we introduced the “Reasoning-to-Coding of Thoughts (RCoT)” approach, which requires the LLM to perform mathematical reasoning in the first turn without any coding instruction, then writing code based on that reasoning. Compared to PoT, RCoT shifts the code distribution towards mathematics in the first turn, prompting more reasoning steps.
Figure 2: RCoT significantly improves the efficiency and effectiveness of the solution. It indicates that our RCoT proves to be more effective, suggesting that it encourages the model to reason critically and find more efficient solutions.
We conducted a comprehensive study with 8 LLMs. Some of our key findings are summarized as follows: (1) The best model, GPT-4o, only solves 26.93% problem in our benchmark, demonstrate the difficulty of our benchmarks. (2) Modern LLMs perform poorly in Graph Theory, Group Theory, Geometry and Topology (Fig. 4). (2) With RCoT, 7 of 8 evaluated LLMs generated more efficient solutions, with most models achieving higher scores (Fig. 2). (3) RCoT can significantly improve the pass@k performance of LLMs (§ 5.4). (4) The quality of reasoning significantly impacts the accuracy and efficiency of the model’s final solution (§ 5.5). More interesting findings can be found in § 5. We hope our findings contribute to a deeper understanding of current reasoning ability of LLMs and the further development of models.
2 Related Work
2.1 Benchmarks
With the rapid development of LLMs, evaluating and exploring the intelligence and limitations of these models has emerged as an urgent issue to address  Chang et al. (2024). Reasoning ability, as a crucial component of general intelligence, has garnered widespread attention since the advent of LLMs  Patel et al. (2021); Cobbe et al. (2021); Valmeekam et al. (2022); Perez et al. (2022); Gupta et al. (2022); Shakarian et al. (2023). Mathematical reasoning, due to its complex mathematical characteristics and rigorous logical relationships, is considered an abstract and high-difficulty task, playing a pivotal role in demonstrating a model’s reasoning capabilities.
To this end, researchers have proposed various benchmarks focused on mathematical reasoning. A natural and mainstream approach is to evaluate LLMs as humans would take math exams, using human exam questions to test their reasoning abilities, categorized by required knowledge levels. Examples include GSM8K at elementary school level, Math and GaokaoBench-Math Zhang et al. (2023) at high school level , College Math  Tang et al. (2024), TheoremQA  Chen et al. (2023), ARB  Sawada et al. (2023) at university level, and OlympiadBench  He et al. (2024), AGIeval-Math  Zhong et al. (2023) at competition level.
Besides, researchers have also introduced many others focused on evaluating various aspects of LLMs like the robustness. These include GSM8K-based variants: GSM-8K-Adv  Anantheswaran et al. (2024), GSM-Hard  Gao et al. (2023), GSM-Plus  Li et al. (2024a), GSM-IC  Shi et al. (2023), and several independent benchmarks: Mathattack  Zhou et al. (2024b), MetaMath  Yu et al. (2023), PROBLEMATHIC  Anantheswaran et al. (2024), MATHCHECK  Zhou et al. (2024a), as well as other benchmarks  Li et al. (2024b, 2023).
The distinctions between our proposed benchmark and existing ones are as follows:
$\bullet$ Multiple Case Validation Instead of using single cases that can be memorized, our questions are sequence-based, allowing numerous cases for validating true understanding.
$\bullet$ True Reasoning Evaluation Hard cases and runtime metrics help filter memorization and compare solution efficiency, precisely assessing reasoning abilities.
2.2 Building Methods
Constructing effective, high-quality datasets is a complex and labor-intensive process. The advent of LLMs offers an opportunity to change this scenario  Valmeekam et al. (2022); Drori et al. (2023); Perez et al. (2022); Chiang and Lee (2023); Liu et al. (2023); Fu et al. (2023); Kocmi and Federmann (2023); Li et al. (2024b). For instance,  Almoubayyed et al. (2023) employed GPT-4o to rewrite mathematics problems based on MATHia  Ritter et al. (2007) to aid students in improving their math performance, validating this with 12,374 students and demonstrating the effectiveness of using LLMs for data construction. These efforts provide a reliable foundation for utilizing LLMs in data processing.
In our study, we utilized GPT-4o to help us deal with data, such as by providing necessary background knowledge for questions and making them more understandable, with more information about the prompts used shown in the appendixC. Subsequently, human verification was performed to ensure consistency before and after LLM usage.
2.3 Prompting Methods
Considering the attributes of large models, they exhibit significant sensitivity to prompts, rendering prompt engineering a critical area of study.
The Chain-of-Thought Wei et al. (2022) prompting technique encourages models to express reasoning steps in natural language before concluding. Similarly, the approach by  Kojima et al. (2022) uses the phrase "Let’s think step by step" to effectively guide large language models through their reasoning.
Inspired by CoT, several effective prompting methods have been developed, such as Tree-of-Thoughts Yao et al. (2024), Graph-of-Thoughts Besta et al. (2024).
Program-of-Thought prompting Chen et al. (2022): PoT generates programs as the intermediate steps and integrates external tools like a Python interpreter for precise calculations, as well as other prompting methods Wang et al. (2023); Gao et al. (2023).
Our Reasoning-Coding of Thoughts (RCoT) method stands out by dividing reasoning into two steps: reasoning and implementing based on reasoning. This segmentation provides deeper insights into LLM reasoning abilities. Coding reduces errors from limited computational skills and evaluates solutions for better success and efficiency. The advantages can be summarized as follows:
$\bullet$ Code Output Evaluation We require LLMs to output code, focusing on reasoning rather than direct answers, to better reflect their reasoning skills.
$\bullet$ Observation of Reasoning Process By mandating code implementation, we can observe and validate the LLM’s reasoning process, not just the final answer.
3 UTMath Benchmark
3.1 Introduction for OEIS.
The OEIS was established to document integer sequences of interest to both professional and amateur mathematicians, and it has become widely cited in the mathematical community. As of February 2024, it contains over 370,000 sequences OEIS Foundation Inc. (2024). Each sequence is accompanied by an identification number, a brief description, some sample integers, generation rules, links to relevant literature, and, where possible, program code for computing the sequences. An example sequence is shown in Apendix A.
3.2 Benchmark Construction
The sequences that meet our expectations and can be used to test reasoning abilities should satisfy the following conditions:
  1. 1.
    The sequence should require reasoning to derive its recurrence or formula, not just simple calculations.
  2. 2.
    The sequence must be derivable with infinitely many terms, allowing deduction of a(n) for any positive integer n.
  3. 3.
    The sequence must be accurately described to deduce it entirely from the description.
  4. 4.
    The sequence should belong to mathematics, preventing reasoning errors from lack of domain knowledge.
Data Crawling
OEIS provides users with a list of principal sequences  [^1] https://oeis.org/wiki/Index_to_OEIS, which are representative sequences in the OEIS. To help users quickly find the main sequences within their areas of interest, OEIS categorizes these sequences into 118 sections based on the first 2-3 letters of their content themes. By scraping the category tags within each section and the AIDs of their subordinate sequences, we obtained 569 categories and 23,238 principal sequences’ AIDs. OEIS provides an interface to request the JSON data of the HTML page for each sequence using its AID  [^2] https://oeis.org/wiki/JSON_Format. By passing the sequence AIDs to this interface, we acquired the JSON data for these 23,238 sequences.
Data Cleaning
We found that some of the sequences we collected did not meet our criteria and should be removed. Here, we present only a portion, with further details provided in the appendix  B.
$\bullet$ Hard to solve, few terms are discoverable A portion of the sequences retrieved are marked as “hard” in the keyword field of their entries in OEIS. According to OEIS, “Any sequence which can be extended only by new ideas, rather than more computation deserves keyword: hard. Similarly, if computing a term of the sequence would probably merit a paper in a peer-reviewed journal (discussing the result, the algorithm, etc.)”  [^3] https://oeis.org/wiki/User:Charles_R_Greathouse_IV/Keywords/difficulty Another related keyword attribute is "fin" (finite), indicating sequences with limited length. For our purposes, an "appropriate" sequence should possess an infinite number of terms and be infinitely derivable.
$\bullet$ Difficult to Generate Programmatically In OEIS, most sequences are provided with fields such as Mathematica, program, or formula, but not all sequences include these fields. Since OEIS does not specify the reasons for missing these fields, we assume that these sequences might be difficult to generate programmatically.
$\bullet$ Overly Simple Some sequences are simple, needing little inference with clear solving algorithms. Due to uncertainty in sequence description, we use GPT-4o to determine if a sequence requires reasoning or just implementation; if mostly implementation, it’s excluded. For example, A000178444https://oeis.org/A000178: "Superfactorials: product of first n factorials.", which is a sequence that merely requires implementation will be removed.
After addressing the aforementioned issues, we ultimately obtained 1053 sequences that meet our criteria.
Standardization of Question Statements
As OEIS is a specialized academic tool, we cannot directly use the descriptions provided for each sequence as problem statements. This is primarily due to the following reasons:
$\bullet$ Specialized Terminology Some sequence descriptions use complex math terms that need examples or explanations to be clear. Using them directly as problems might test mathematical knowledge rather than reasoning skills. So, it’s important to explain key concepts to focus on reasoning and reduce the extra knowledge needed.
$\bullet$ Brevity and Ambiguity Some sequence descriptions are excessively brief and lack a clear definition of what a(n) is.
We used GPT-4o to standardize these by adding background info without revealing solutions and making the language smoother. More details are in the appendix 10. Despite our requirement that ensuring the consistency of meaning between the original and processed descriptions, hallucinations can still occur. To mitigate this, we performed a manual verification of the standardized problem statements to ensure they matched the original descriptions in meaning and were easy to comprehend.
Figure 3: Comparison between original and standardized problem statements. The standardized version includes hints and explains the specific meaning of $\sigma(x)$.
Hard Test Cases Mining
The primary goal of this paper is to evaluate the reasoning ability of LLMs. Generally, more efficient solutions to a problem imply stronger reasoning capabilities. Therefore, we aim for our evaluation to distinguish whether a solution is efficient. However, in the OEIS (Online Encyclopedia of Integer Sequences), each sequence only lists the first few n terms, normally n<100, which can be obtained without requiring particularly efficient methods. This limitation prevents the evaluation from effectively distinguishing between efficient and inefficient solutions. An obvious fact is that the difficulty of computing the first 10 terms of a sequence within a time limit is significantly different from computing terms starting from the $10^{6}$th term. Therefore, we aim to create more challenging test data to better assess the reasoning capabilities of LLMs.
Fortunately, many OEIS sequences include corresponding Mathematica code that can be regarded as the ground-truth solution for each problem. We extract this Mathematica code for each sequence, formalizing it to compute the first $N$ terms, $A_{1},...,A_{N}$, of the sequence. Using a binary search algorithm, we determine the maximum value of $N_{max}$ for which the code can compute the sequence within 10 seconds, where we set $10^{6}$ as the upper bound. Finally, we add the last 10 terms $A_{N_{max-9}},...,A_{N_{max}}$ into our benchmark as the hard test cases to evaluate the complexity of a solution. Our experiments demonstrate that these cases precisely differentiate more efficient and intelligent solutions.
3.3 Evaluation Metrics
We adopt the metric pass@$k$ to evaluate the performance of LLMs. The metric pass@$k$ is a classic metric in code generation, where k code samples are generated per problem, a problem is considered solved if any sample passes the unit tests, and the total fraction of problems solved is reported. We use the stable method of calculation proposed by (Chen et al., 2021):
[…]
3.4 Dataset Statistics
To gain a deeper understanding of the composition of the UTMath Benchmark, we analyzed the potential fields each problem might pertain to and identified nine mathematical fields: Number Theory, Graph Theory, Group Theory, Game Theory, Discrete Mathematics, Combinatorial Mathematics, Geometry and Topology, Polynomial and Series Expansions, Special Numbers, Formal Languages. Using GPT-4o, we allocated each problem to these fields and obtained the distribution shown in Table 1.
Table 1: Categories and distribution of problems
Category Number
Number Theory 159
Graph Theory 79
Group Theory 65
Discrete Mathematics 158
Combinatorial Mathematics 158
Geometry and Topology 70
Poly. and Series Expan. 151
Special Numbers 157
Formal Languages 56
Total Problems 1053
Table 2: Pass Rate and Average Run Time of LLMs on UTMath
Model Pass@1 (%) $\uparrow$ Pass@5(%) $\uparrow$ Avg. Run Time (s) $\downarrow$
PoT RCoT PoT RCoT PoT RCoT Efficiency
closed-source models
GPT-4o 25.53 26.93 ($+$1.40) 32.67 35.90 ($+$3.23) 6.98 6.23 $+$12.04%
Claude-3.5-Sonnet 18.58 19.11 ($+$0.53) 27.83 31.34 ($+$3.51) 6.44 5.32 $+$21.05%
Gemini-1.5-Pro 14.13 12.35 (--1.78) 25.17 24.69 (--0.48) 5.71 5.76 --0.87%
GPT-3.5-Turbo 11.68 6.82 (--4.86) 17.09 13.30 (--3.79) 5.42 5.06 $+$7.11%
open-source models
Qwen2.5-72B 23.48 22.17 (--1.31) 31.05 33.33 ($+$2.28) 5.88 4.31 $+$36.42%
DeepSeek-V2.5-236B 20.95 21.63 ($+$0.68) 30.10 31.72 ($+$1.62) 6.64 5.44 $+$22.06%
Qwen2.5-Math-72B 19.72 20.53 ($+$0.81) 26.69 28.11 ($+$1.42) 5.04 3.81 $+$24.40%
LLaMA-3.1-405B 15.76 16.09 ($+$0.33) 25.26 27.35 ($+$2.09) 5.73 5.12 $+$11.91%
4 Reasoning-to-Coding of Thoughts
Compared to evaluation methods that solely rely on checking if the results generated by LLMs are identical, without assessing the reasoning process, we propose a method based on evaluating mathematical reasoning processes through code implementation.
Initially, we adopted the Program of Thought (PoT) method, where the LLM had to perform reasoning and implement it in one step. However, we noticed that the LLMs often resorted to simpler algorithms, which led to high time complexity or even failure with more complex problems due to limited reasoning depth. To improve this, we explored the Reasoning-to-Coding of Thoughts framework, which separates reasoning and implementation into different steps.
In the first round, LLM focuses only on reasoning about the problem. Compared to PoT, where reasoning is described broadly, dedicating an entire round to reasoning allows the LLM to generate a step-by-step, detailed logical reasoning chain, including mathematical theorems, formulas, and properties used. This deeper reasoning approach reduces the likelihood of errors and facilitates the creation of less complex algorithms.
In the second round, the LLM implements the reasoning process generated in the first round.This gives us new perspectives beyond the final answer when evaluating reasoning abilities. By converting the computation process into code, we avoid errors caused by low computational skills, allowing us to gain more accurate and genuine insights into the LLM’s reasoning abilities. Additionally, assessing the pass rate and runtime of the code provides a method to evaluate reasoning strategies, where better approaches should have higher accuracy and lower time complexity.
By dividing the process into two steps, The detailed reasoning in the first round ensures that the code implementation in the second round closely aligns with the reasoning process, reducing mismatches between reasoning and implementation.
5 Experiment
Figure 4: Performance comparison of models across PoT and RCoT tasks at different pass@k levels.
5.1 Experimental Setup
Here, we consider the closed-source models, i.e., GPT-3.5-Turbo/GPT-4o from OpenAI OpenAI (2024), Claude-3.5-Sonnet Claude (2024), Gemini-1.5-Pro Reid et al. (2024), as well as the open-source models, i.e., LLaMA-3.1 Dubey et al. (2024), Qwen2.5 Qwen (2024a), Qwen2.5-Math Qwen (2024b), DeepSeek-V2.5 Bi et al. (2024). The metric pass@$1$ is calculated as the average result over 5 run times. We run all evaluations in a laptop with CPU Intel(R) Core(TM) i7-10750H CPU @ 2.60GHz.
5.2 Evaluation on UTMath
Here we evaluate both open-source and closed-source models using RCoT and PoT in Tab. 2. The experimental results shows that all tested models performed poorly on our benchmark. The best model, GPT-4o, only solves 26.93% problem in our benchmark. Since our problems are sourced from the OEIS, they consist of sequences and solutions proposed by various mathematicians in the context of cutting-edge research. This suggests that our benchmark is challenging enough to help guide future directions for improving LLMs.
Compared to PoT, our method RCoT demonstrates superiority in two aspects. First, prompting with RCoT achieves higher pass@$5$ performance across 6 LLMs, with the best results observed on GPT-4o. Second, the solutions generated by RCoT demonstrate more efficient performance, particularly Qwen2.5-72B, where the RCoT approach achieves an efficiency improvement of over 36.42% compared to PoT, as shown in Tab. 2 and Fig. 2. It indicates that, with RCoT prompting, the model engages in deeper reasoning, significantly reducing solution complexity and enhancing solution performance. However, some models experienced a decrease in pass@$1$ with RCoT . Gemini-1.5-Pro, GPT-3.5-Turbo, and Qwen2.5-72B’s accuracies slightly dropped. Qwen2.5-72B’s dropped on pass@$1$ but increased on pass@$5$. It indicates that RCoT brings more room in multiple inference times. The poorer performance may be due to the fact that attempting to devise a more efficient solution often involves a higher level of difficulty, and these models are more prone to errors when reasoning through more complex solutions.
5.3 The Effectiveness of Hard Test Cases
As we mentioned in Sec. 3.2, each sequence in the OEIS lists only the initial terms, which we refer to as “easy test cases”. To investigate the model’s ability to handle challenging cases, we evaluated whether it could predict values that appear later (i.e., $10^{6}$) in a sequence. These later values are typically underrepresented in pre-training data and often require more computation time and a more precise implementation to retrieve accurately. The experimental results, which depicted in Tab. 3, reveal that the model’s performance drops significantly when handling these hard cases. This indicates that introducing these cases can prevent simple solutions from passing all the easy cases, thereby filtering for more advanced solutions. Our benchmark provides a better measure of the reasoning abilities across different models.
Table 3: Performance of different models on easy and hard test cases. Divide into open-source and closed-source models, and sort each model in descending order based on the performance on Easy & Hard cases.
Model Easy Easy & Hard
GPT-4o 34.95 26.93
Claude-3.5-Sonnet 24.86 19.11
Gemini-1.5-Pro 16.35 12.35
GPT-3.5-Turbo 8.72 6.82
Qwen2.5-72B 28.96 22.17
DeepSeek-V2.5 27.52 21.63
Qwen2.5-Math-72B 24.60 20.53
LLaMA-3.1-405B 22.09 16.09
Table 4: Performance on Different Problem Categories.(%) Categories are represented by abbreviations. NT: Number Theory; T.: Theory; DM: Discrete Mathematics; CM: Combinatorial Mathematics; GT: Geometry and Topology; PSE: Polynomial and Series Expansions; SN: Special Numbers; FL: Formal Languages;
Model NT Graph T. Group T. DM CM GT PSE SN FL pass@1
closed-source models
GPT-4o 43.9 2.78 11.69 38.23 24.94 3.43 16.42 33.89 42.5 26.93
Claude-3.5-Sonnet 33.58 1.52 8.0 29.49 12.91 5.43 11.52 26.62 20.36 19.11
Gemini-1.5-Pro 22.64 1.27 5.23 19.75 8.35 2.29 8.61 14.01 15.36 12.35
GPT-3.5-Turbo 13.08 0.0 1.85 11.39 3.29 0.29 2.78 10.96 8.93 6.82
open source models
Qwen2.5-72B 36.86 2.53 12.0 30.63 15.95 6.0 18.15 29.43 24.29 22.17
DeepSeek-V2.5 38.24 1.27 8.92 33.16 17.34 2.29 12.45 31.08 20.0 21.63
Qwen2.5-Math-72B 35.35 1.27 8.62 28.73 14.81 4.0 17.48 28.15 20.0 20.53
LLaMA-3.1-405B 29.56 0.76 4.92 25.44 9.62 2.0 9.54 22.55 21.43 16.09
5.4 Scaling of the Inference Times
We compared the performance difference between running the LLMs five times and reported the metric of pass@$k$. As shown in Fig. 4, all models improved their performance with an increasing number of inference times. For Qwen2.5-72B, RCoT was slightly weaker than PoT in the first inference but quickly approached and surpassed PoT in subsequent run times. For GPT-3.5-Turbo and Gemini-1.5-Pro, PoT consistently outperformed RCoT. We observed that with an increasing number of inference time, RCoT consistently demonstrated a growing advantage in performance across almost all models, except for GPT-3.5. This suggests that RCoT may perform better in models with stronger reasoning capabilities.
Figure 5: Performance comparison between self-reasoning and using GPT-4o reasoning for coding across different models. The results show that models perform better when relying on GPT-4o’s reasoning output.
5.5 Importance of the Reasoning Step
GPT-4o has the best performance, while we are unclear whether this was due to its superior reasoning or coding ability. To investigate further, we tested using GPT-4o for the reasoning step while other models perform the coding step based on the reasoning result from GPT-4o. As depicted in Fig.5, the results showed that the performance of models, except for Qwen2.5-72B, increased significantly when implementing coding based on GPT-4o’s reasoning output, suggesting that the reasoning quality is important and GPT-4o does produce higher-quality reasoning results.
5.6 Performance on Different Categories
Our benchmark comprehensively evaluates the LLMs’ ability across various categories of math problems. We analyzed the differences in performance in terms of categories. GPT-4o continued to demonstrate a clear advantage across most tasks, achieving the best performance in five domains. Notably, it outperformed the second-best model by 7.6% in combinatorial mathematics and by 18.21% in formal languages. In contrast, as an open-source model, Qwen2.5-72B secured the top spot in the other three domains. All models performed very poorly in the categories of Graph Theory, Group Theory, and Geometry and Topology, with accuracy rates below 12%, highlighting the need for further exploration in these areas.
Appendix A An Example Sequence in OEIS
An example sequence is shown in Fig. 6
Figure 6: Sequence A001776 in OEIS.
Appendix B Rulers for Data Cleaning
  1. 1.
    Problem Description: The sequence is too difficult, requiring extensive background knowledge, or only a limited number of terms are found.
    Filtering Method: Remove sequences with keywords containing ‘hard’, ‘fin’ (finite).
  2. 2.
    Problem Description: The sequence is hard to generate with a program.
    Filtering Method: Check if it contains program, formula, or Mathematica fields in the sequence’s json data.
  3. 3.
    Problem Description: The sequence is too simple with an explicit recurrence or closed formula.
    Filtering Method: Search if the description includes ‘a(n) =’.
  4. 4.
    Problem Description: Solving the sequence requires information from other OEIS sequences.
    Filtering Method: Search if the sequence’s description contains the AID of other sequences (‘A’ + six-digit number with leading zeros).
  5. 5.
    Problem Description: The sequence is decimal expansion of a certain number.
    Filtering Method: Search if the description includes ‘decimal’.
  6. 6.
    Problem Description: The sequence consists of repetitions or a constant value.
    Filtering Method: Search if the description includes both ’repeat’ and ’period’ or ’constant sequence’.
  7. 7.
    Problem Description: The description is too vague.
    Filtering Method: Search if the description includes ‘related to’.
  8. 8.
    Problem Description: Another version of a concept.
    Filtering Method: Search if the title includes ‘another version’, ‘second kind’, etc.
  9. 9.
    Problem Description: The sequence is formed by taking mod of a constant.
    Filtering Method: Search if the description includes ‘module’.
  10. 10.
    Problem Description: The values in the sequence are too large, which might cause LLM tokenization errors.
    Filtering Method: Check if any term’s length exceeds 18 digits(i.e., greater than $181e$ 18), remove it.
  11. 11.
    Problem Description: Coefficient triangles or ‘read by row’ topics.
    Filtering Method: Search if the sequence’s description includes ‘read by row’, ‘triangle of coefficient’.
  12. 12.
    Problem Description: The description is too short, either purely implementation or lacks necessary information.
    Filtering Method: Check if the title length is below 5.
  13. 13.
    Problem Description: More like a reasoning puzzle.
    Filtering Method: Use GPT-4o to judge, with the prompt outlined in the prompts section of the appendix 9.
  14. 14.
    Problem Description: Non-mathematical topics.
    Filtering Method: Use GPT-4o to judge, with the prompt outlined in the prompts section of the appendix 11.
Appendix C Prompts
Figure 7: The prompts for Program-of-Thoughts
Figure 8: The prompt for Reasoning-to-Coding of Thoughts.
Figure 9: The prompt used to determine reasoning / implementation questions.
Figure 10: The prompt used to standardize question statements.
Figure 11: The prompt used to determine whether the question belongs to the field of mathematics.